OpenStudy (anonymous):

Kobe's overtime pay is $5 an hour more than his regular pay. He worked 8 hours at his regular wage and 3 hours at his overtime wage. He earned $114. What is Kobe's regular wage per hour? $5 per hour $9 per hour $14 per hour $23 per hour

OpenStudy (anonymous):

8x+3(x+5)=114, 8x+3x+15=114, 11x=99, x=99/11, x=9. His regular wage per hour is $9.00.

OpenStudy (anonymous):

Thank you.